OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4-x and y = 2x + 3 intersect are the solutions of the equation 4-x = 2x + 3. Part B: Make tables to find the solution to 4-x = 2x + 3. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4-x = 2x + 3 graphically?

OpenStudy (jim766):

A. if the points are the same, the y's will be the same. y = 4-x and y = 2x + 3 , they both equal y, so you can substitute one in for the other. y = 4-x , y = 2x + 3 4-x = 2x + 3

OpenStudy (anonymous):

I'm sorry I typed the questions wrong. It should look like this: Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4^-x and y = 2^(x + 3) intersect are the solutions of the equation 4^-x = 2^(x + 3). Part B: Make tables to find the solution to 4^-x = 2^(x + 3). Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4^-x = 2^(x + 3) graphically?

OpenStudy (jim766):

same logic since the both equal y, they must equal each other

OpenStudy (jim766):

table for 4^-x -3 64 -2 16 -1 4 0 1 1 .25 2 .065 3 .01563

OpenStudy (anonymous):

Shouldn't the y values on that table be switched?